OpenStudy (anonymous):

soybean meal is 16% protein; cornmeal is 8% protein how many pounds of each should be mixed together in order to get 320lb mixture that is 10% protein

OpenStudy (anonymous):

i think i have done this exact problem maybe 8 times. write \[.16x+.08(230-x)=32\] and sovle for x

OpenStudy (anonymous):

actually it is \[.16x+.08(320-x)=32\]

OpenStudy (anonymous):

Use x + y = 320 and .16x + .08y = .10 Solve the system

OpenStudy (anonymous):

i get 80 of course you can use two varialbes if you like

OpenStudy (anonymous):

but if you do it should be \[.16x+.08y=32\]

OpenStudy (anonymous):

\[.16x+.08(320-x)=32\] because if you use x soybeans you use 320-x cormeal and 10% of 320 is 32

OpenStudy (anonymous):

so how many pounds of cornmeal should be in the mix and how many pounds of soybean should be in the mix

OpenStudy (anonymous):

\[.16x+25.6-.08x=32\] \[.08x+25.6=32\] \[.08x=32-25.6=6.4\] \[x=\frac{6.4}{.08}=80\]

OpenStudy (anonymous):

80 cornmeal the rest soybeans

OpenStudy (anonymous):

I am subtracting 32 from 80 to get the soybean

OpenStudy (anonymous):

or is the soybean 32

OpenStudy (anonymous):

Its a lot easier to me if you use the substitution method

OpenStudy (anonymous):

tervin03 can you show me how to arrive at the answer by using that method

OpenStudy (anonymous):

0Use x + y = 320 and .16x + .08y =32 and use the substitution method. So in the first equation x + y = 320 you can subtract x from both side to get y = 320 - x and substitute it in the second equation making it .16x +.08(320 - x) =32

OpenStudy (anonymous):

Solve it for x and then back substitute the answer in the first original equation x + y = 320 to get y

OpenStudy (anonymous):

Tervin03 so how many pounds of cornmeal should be in the mix and how many pounds of soybean should be in the mix

OpenStudy (anonymous):

Soybean is 80 and cornmeal is 240